OpenStudy (anonymous):

Find the y-intercept and x-intercept of the following equation: -9x-6y=-54. What are the ordered pairs for each?

OpenStudy (anonymous):

solve for each variable, then plug in 0 for that variable and solve for the other. y = 9+(9/6)x, 0 = 9+(9/6)x, and solve for x, this would be (x,0), whatever x is, x intercept do the same thing for x

OpenStudy (anonymous):

you want to solve for one variable, meaning youre looking for, for exmple, y= something. to do that, you need to cancel out everything thats on the y's side except for the y and we can do this by inverse operation. inverse of - is + and inverse of * is /, and visa versa. to remove -9x, we can add to the equation +9x. and remember to do that on both sides of the equation, we we have -6y = -54 + 9x. now, to get rid of -6, remember that -6y is -6 * y, so we divide it by -6, to get y = (-54 + 9x)/-6, and simply to y = 9 + (3/2)x

OpenStudy (anonymous):

9/6 is equivalent to 3/2

OpenStudy (anonymous):

Sorry it looks like I made it more complicated, it's not necessary to solve for each variable, you can plug in 0 for y and solve for x from there which will give you x-intercept in (x,0) and 0 for x and solve for y from there which will give you y-intercept in (0,y)

OpenStudy (anonymous):

-9x-6y=-54 x = 0, y = 54/ 6 = 9 ( 0, 9) y = 0 , x = 54/9 = 6 ( 6, 0)